# Loki: Low-Rank Keys for Efficient Sparse Attention

### v1

@@ Abstract
Inference on large language models can be expensive in terms of the compute and memory costs involved, especially when long sequence lengths are used. In particular, the self-attention mechanism used in such models contributes significantly to these costs, which has resulted in several recent works that propose sparse attention approximations for inference. In this work, we propose to approximate the self-attention computation by focusing on the dimensionality of key vectors computed in the attention block. Our analysis reveals that the key vectors lie in a significantly lower-dimensional space, consistently across several datasets and models. Exploiting this observation, we propose Loki, a novel sparse attention method that ranks and selects tokens in the KV-cache based on attention scores computed in low-dimensional space. Our evaluations show that Loki is able to maintain the efficacy of the models better than other popular approximation methods, while speeding up the attention computation due to reduced data movement (load/store) and compute costs.
@@ 1 Introduction
As the sizes of large language models (LLMs) increase, deploying them for efficient inference remains a significant challenge, primarily due to the computation and memory access bottlenecks in the self-attention block [31]. These challenges arise due to the auto-regressive nature of attention, where the output is generated token by token. At each generation step, the entire preceding state, represented by the key-value (KV) cache [23], must be fetched from memory. This state can, at times, be larger than the model parameters themselves [17]. The cost of reading the KV cache from the GPU DRAM to the registers at each generation step is prohibitively expensive. During inference, this KV-cache read/write cost scales quadratically with the sequence length, as opposed to training where it scales linearly. Moreover, the computational cost of matrix multiplications in the attention layers also scales quadratically with the sequence length [36].
@@
Several strategies [38, 25, 18] have been proposed to address this challenge by reducing the computational complexity and/or memory demands associated with the self-attention mechanism. One promising category of approaches focuses on approximating attention, employing techniques such as quantization (compression along the bit dimension) or using a subset of the the number of tokens in the KV cache [9] (sparse attention).
While other sparse attention methods either permanently throwing away tokens [38] or have a fixed sparsity pattern [33], we focus on using the keys (in the feature dimension) to decide which tokens might be important. We hypothesize that the keys in the KV cache lie in a low-dimensional space. We begin with exploring the intrinsic dimensionality of the key vectors in the KV cache. We perform principal component analysis (PCA) [15] on the key vectors within each layer and head using a calibration dataset to study their dimensionality. As shown in Figure 1 (left), we find that the key vectors indeed lie in a low-dimensional space for several popular and recent models like Llama-3 [2] and Mixtral [14].
Next, we exploit the low dimensionality of the key vectors to develop a sparse attention method that does not significantly sacrifice model quality and also reduces data movement and compute costs. We call our method ‘Loki’. We store all the principal components of the PCA-transformed keys, but use the first to compute approximate attention scores for picking the top- tokens. We then use the dimensionality only for the selected keys to compute the final attention scores. Figure 1 (right) provides an overview of Loki.
@@
Our theoretical complexity analysis shows that Loki can provide significant speedups in the attention step. However, empirically, this requires efficient implementation of Loki to minimize data movement in the additional operations introduced. Thus, we implement optimized matrix multiplication kernels for Loki, leading to a speedup of up to 40% over base attention for the Llama2-13B model. For this speedup setting, the average degradation in model accuracy (measured across 6 different benchmarks and 8 different models) is only 6.8% points.
@@
Loki: a sparse attention method that exploits the aforementioned low dimensionality of keys to speedup attention computation without sacrificing model quality.
@@
Evaluation of Loki on multiple LLMs and downstream tasks, showing that it can achieve significant speedups with minimal degradation in model quality.
@@ 2 Background and Related Work
where is the query, and and are the key and value caches respectively. Additionally, newer transformer models add Rotational Positional Embeddings (RoPE) [28] to the keys and query, before computing the attention scores. Since every query attends to all past keys, the attention mechanism has a quadratic complexity in the number of tokens.
@@ 2.1 Related Work
Numerous studies have explored the low-rank structures in transformers for various purposes. Linformer [32] demonstrated that the attention score matrix is low-rank and proposed alternative attention formulations through low-rank factorizations during training for linear computational complexity. LoRA [11] showed that the weights of transformers and the updates to these weights during fine-tuning reside in a low-dimensional subspace. To the best of our knowledge, our work is the first to study the intrinsic low dimensionality of the attention keys themselves and demonstrate the generalizability of this low-dimensional structure in a post-training setting (for natural language data).
@@
Sparse-transformers [5] was one of the first works to introduce a sparse-attention method employing fixed or strided sparsity patterns in the attention mechanism. Reformer [16] used locally-sensitive hashing to compute attention scores in a sparse manner. Performer [6] used positive orthogonal random features to approximate the attention mechanism. Unlike these methods, which require training or fine-tuning, our approach operates entirely post-training without any fine-tuning.
@@
Another category of sparse attention methods employ token eviction policies to permanently delete tokens from the KV-cache based on some heuristic. StreamingLLM [33] introduced the concept of attention sink, using initial tokens and a rolling KV-cache for processing infinite-length sequences. Zhang et al. [38] retain only "Heavy Hitters" tokens in the KV-cache based on accumulated attention scores. Liu et al. [18] propose Scissorhands, which prioritizes important tokens based on the "Persistence of Importance Hypothesis". Ge et al. [8] propose an adaptive eviction policy for each transformer layer. These methods are effective in reducing the memory and compute footprint of the attention but suffer from permanent loss of information that can lead to generation a non-trivial degradation in model quality. Our method does not involve any permanent loss of information with the trade-off of not reducing the memory footprint. Quantization-based approximate approaches [12, 21] are complimentary to our work and can be applied in tandem.
@@
SPAR-Q Attention [25] is a recent work that inspires our approach. They use the maximum magnitude dimensions of the queries, and the corresponding dimensions of the keys to compute approximate attention scores, and then use the full attention scores for the top-k keys. However, the main drawback of their method is that it requires expensive non-contiguous indexing of columns of the keys matrix. Further, they store two copies of the past keys leading to a 50% increase in memory usage. Loki does not require additional memory, and the natural ordering of the principal components allows one of the two indexing operations to be replaced with a more efficient slicing operation.
@@ 3 Dimensionality Analysis of Attention Keys
In this section, we analyze whether most of the variance in the keys can be explained using dimensions , where is the full dimensionality of the keys.
3.1 Setup
To investigate the dimensionality of attention keys, we run 11 transformer-based models: Llama-2 7B/13B/70B [30], Llama-3 8B/70B [2], TinyLlama-1.1B [37], Pythia-6.9B [4], Mistral-7B [13], Mixtral-8x7B/8x22B [14], and Phi3-Mini-4K [20] on 3 popular English language datasets datasets: WikiText-2 [19] (Validation Split), C4 [24] (Custom Split), and BookCorpus [39] (Custom Split). Custom splits are used for datasets where the validation split is not available. We run perplexity evaluation on these datasets and save the generated attention keys, before and after the application of rotary embeddings [28], referred to as pre-rotary and post-rotary keys, respectively throughout the paper. We then perform PCA on all the keys generated for each layer and head individually.
@@ 3.1 Setup
where, is the normalized eigenvalue of the covariance matrix of the keys for the layer and head. We average this metric ranks across all heads of the layer and refer to it as .
3.2 Analysis and Discussion
Figure 1 (left) shows the average averaged across all layers for models with full key dimensionality of 128. We can see that the average rank is significantly lower than the full dimensionality of the keys for all models. Diving deeper, we present a layer-wise analysis for a few models: Llama2-7B, Llama3-70B, Mixtral-8x7B, and Phi3-Mini-4K in Figure 2. The results for the other models are similar and can be found in the appendix A.1.
We observe that the dimensionality of the keys (both pre-rotary and post-rotary) is significantly lower than the full dimensionality of the keys across all calibration datasets. Furthermore, the for a particular layer is consistent across datasets, for all combinations of models and datasets. This indicates that the lower-dimensional structure of the keys is consistent when calculated using different calibration datasets. Another trend we observe is that the initial layers of most models have a very low rank, as compared to the later layers, and this trend is particularly prominent for the pre-rotary keys . Lastly, we also observe that for most models, the average of across all layers is lower for pre-rotary keys as compared to post-rotary keys, indicating that the rotary embeddings increase the dimensionality of the keys. Further analysis on the variation of the rank across different heads within a layer and across different layers within a model can be found in the appendix.
@@ 3.2 Analysis and Discussion
4 Method: Loki
In this section, we present Loki, a novel sparse attention method leveraging PCA for top- attention computation (i.e. computing attention with a sized subset of the KV-cache tokens). We first prove some theoretical properties of attention in the PCA-transformed space, and then present the Loki algorithm. We also present a memory-efficient kernel the top- attention computation, crucial for empirical speedups with our method.
4.1 Properties of Attention in the PCA Transformed Space
First, we introduce some lemmas that provide the rationale for our approach to computing attention in the PCA-transformed space.
@@ 4.1 Properties of Attention in the PCA Transformed Space
Let be the PCA projection matrix calibrated offline on a dataset. Assuming we are generating the token in the sequence, let be the query vector for the token, be the key vectors, including the past keys and the current key. Then, the attention scores computed using the PCA-transformed query and keys are equivalent to the attention scores computed using the original query and keys.
@@
Let be an orthogonal transformation that transforms the keys into the reduced dimension space as . Our objective is to minimize the following reconstruction error:
@@
We change our objective to minimize the upper bound on the RHS instead of the original objective. We know that PCA minimizes the reconstruction error (2nd term in the RHS) among all the orthogonal transformations. Thus, it follows that the optimal value of , and ∎
@@
Using the above lemmas and our dimensionality analysis showing that the key vectors have low intrinsic dimensionality, we are now ready to present the algorithm for Loki.
4.2 Algorithm and Complexity Analysis
In this section, we propose the algorithm for our PCA-based Top-K Attention approach (Loki). Previous works [34, 29] have shown that attention scores for a query are highly concentrated on a small subset of keys. This observation has led to several methods to compute attention using only the top- keys. However, these previous works either compute the exact attention scores and then select the top- keys [10] or compute non-exact scores but have significantly higher memory requirements [25]. Loki alleviates these issues by computing approximate attention scores (for ranking the keys) in the reduced lower-dimensional space, without any significant increase in memory requirements. Algorithm 1 shows our Loki method. Line 5 of the algorithm computes the approximate attention scores using principal dimensions of the query and key vectors. Lines 6-7 select the top- keys based on the approximate attention scores. Line 8 computes the exact attention scores using the selected top- keys, directly in the transformed space (Lemma 4.1).
Compute and Memory Analysis: For vanilla attention, the complexity of computing is and the complexity of multiplying the values with the attention scores is . For Loki, the complexity of calculating the approximate attention scores (Line 5) is . The complexity of selecting the top-K keys (Lines 6-7) is approximately . The complexity of calculating the exact attention scores and multiplying with the values (Line 8-9) is . Additionally, the complexity of projections into the PCA space (Line 3) is . Assuming the complexity of selecting the top- keys is small compared to the other operations, the overall complexity of the algorithm is . Then, we have:
@@ 4.3 Implementation in Triton
Performing Loki efficiently involves complex indexing operations within the KV-Cache (lines 5 and 7 of Algorithm 1). Standard PyTorch operations create temporary, dense copies of the KV-Cache data in memory, leading to slowdowns due to expensive memory access. To alleviate this issue, we develop optimized kernels in Triton [1] for the three matrix multiplication operations in Loki. Our kernels can directly access relevant subsets of the KV-Cache (both feature and sequence dimensions) and perform computations within GPU registers. This eliminates the need for creating dense copies, significantly improving performance. Our approach builds on SPAR-Q [25], which introduced similar kernels for top-k attention calculations. However, we identified and addressed inefficiencies in the SPAR-Q kernels, which resulted in speedups of nearly in certain scenarios. (see Appendix C).
5 Evaluation
In this section, we present our experimental setup and evaluation of Loki via three studies: (1) Comparing Loki with baselines on common ML benchmarks, (2) Examining its generalizability across various calibration datasets, and (3) Benchmarking its computational efficiency against vanilla attention computation.
5.1 Experimental Setup
For our ML benchmarking, we evaluate our method using two approaches - perplexity evaluation on WikiText-2 [19] dataset (test split) and downstream task performance using the LLM-harness benchmark [7]. For the downstream tasks, we choose the same tasks as the HuggingFace OpenLLM leaderboard [3]: The metrics used for these tasks are identical to the leaderboard
We compare our method against 3 baselines - full attention without any approximations, the exact TopK approach which computes the exact attention scores and then uses the top-k tokens to compute the final output, and H2O [38] method which is a popular token-eviction method. For these comparisons, we show the results with a budget size of = 0.25 and 0.125. For our method, we additionally use = 0.25 and 0.125. This configuration of our represents a 2.6x theoretical speedup. Table 1 provides an overview of the baseline methods and the associated budget terms. H2O’s budget was split equally between the heavy hitter and recent tokens, as per the original paper. For H2O, we were unable to run the GSM8K task as the the author’s ML benchmarking code was too memory intensive to run for that task. For the aforementioned experiments, we generate PCA transforms using the WikiText-103 dataset. For the generalizability study, we compare the results of our method with PCA transforms from different calibration datasets: WikiText-103 [19], C4 [24], and BookCorpus [39]. Additionally, we also benchmark our triton based implementation of PCA for running attention in a Llama2-13B-like setup (same hidden size and number of heads) for various prompt and generation lengths, and demonstrate speedups over vanilla attention.

### v2

@@ Abstract
Inference on large language models (LLMs) can be expensive in terms of the compute and memory costs involved, especially when long sequence lengths are used. In particular, the self-attention mechanism used in LLM inference contributes significantly to these costs, which has sparked an interest in approximating the self-attention computation to reduce such costs. In this work, we propose to approximate self-attention by focusing on the dimensionality of key vectors computed in the attention block. Our analysis reveals that key vectors lie in a significantly lower-dimensional space, consistently across several datasets and models. Exploiting this observation, we propose Loki, a novel sparse attention method that ranks and selects tokens in the KV-cache based on attention scores computed in low-dimensional space. Our evaluations show that Loki is able to speed up the attention computation due to reduced data movement (load/store) and compute costs while maintaining the efficacy of the models better than other popular approximation methods.
@@ 1 Introduction
As large language models (LLMs) grow in size, deploying them for efficient inference presents substantial challenges, largely due to computation and memory access bottlenecks in the self-attention block [32], especially when handling long sequences. These challenges stem from the autoregressive nature of attention, which generates the output one token at a time. At each step, the entire preceding state, stored in the key-value (KV) cache, must be fetched from memory, which can sometimes exceed the size of the model parameters itself [18]. This frequent KV-cache access from GPU DRAM to registers becomes costly, as it scales quadratically with the output sequence length. In addition, matrix multiplications in the attention layers also have a quadratic scaling cost with sequence length, compounding the overall computational burden.
@@
Several strategies [39, 26, 20] have been proposed to address this challenge by reducing the computational complexity and/or memory demands associated with the self-attention mechanism. One promising category of approaches focuses on approximating attention, employing techniques such as quantization or using a subset of the tokens in the KV-cache [11] (sparse attention).
In contrast to other sparse attention approaches that either permanently prune tokens from the key-value cache [39] or impose a fixed sparsity pattern [35], our proposed method dynamically selects key tokens at each generation step based on approximate attention scores and avoids deletions. This approach is inspired by a critical observation: across a range of LLMs and datasets, key tensors consistently occupy a significantly lower-dimensional space than the full attention head dimension. For instance, in Figure 1 (left), we show that across various LLMs [8, 16], 90% of the variance explained by PCA is captured at an effective key vector rank of around 80, despite the key tensor dimension being much larger (128).
Based on this observation, we introduce Loki, a sparse attention method that leverages the low-dimensional structure of key vectors to reduce data movement and computation costs without significantly impacting model quality. First, we apply PCA to keys generated from a calibration dataset, storing all principal components but using only the top (25-50%) to compute approximate attention scores during inference. This dimensionality reduction, informed by our previous observation that key vectors have low effective rank, allows us to efficiently identify the top- (12.5-25%) most relevant tokens using the approximate scores. For these selected keys, we then revert to the full dimensionality to compute the final attention scores, ensuring both efficiency and accuracy. Figure 1 (right) illustrates our approach.
@@
Our theoretical complexity analysis demonstrates that Loki can provide significant speedups in the attention step. However, actually realizing these gains requires an efficient implementation of our method to minimize data movement in the additional operations introduced on top of the original self attention algorithm. Thus, we implement optimized sparse matrix multiplication kernels for Loki in Triton, leading to a speedup of up to 45% over the standard HuggingFace Transformer’s [34] attention implementation (vanilla attention) for Llama2-13B. For this setting, the average degradation in model accuracy (measured across 6 different benchmarks and 8 different models) is only 6.8%.
@@
Loki: a sparse attention method that exploits the aforementioned low dimensionality of keys to make the attention computation faster without sacrificing model quality.
@@
Evaluation of Loki 111https://github.com/hpcgroup/loki on multiple LLMs and downstream tasks, showing that it can achieve significant speedups with minimal degradation in model quality.
@@ 2 Background and Related Work
where is the query, and and are the key and value caches respectively. Additionally, newer transformer models add Rotary Position Embeddings (RoPE) [29] to the keys and query, before computing the attention scores. Since every query attends to all past keys, the mechanism has a quadratic complexity in number of input generated tokens.
@@ 2.1 Related Work
Numerous studies have explored the low-rank structures in transformers for various purposes. Linformer [33] demonstrated that the attention score matrix is low-rank and proposed alternative low-rank attention formulations during training for linear computational complexity. LoRA [13] showed that parameter updates to a transformer model during fine-tuning reside in a low-dimensional subspace. To the best of our knowledge, our work is the first to study the intrinsic low dimensionality of the attention keys themselves and demonstrate the generalizability of this low-dimensional structure across different models (for natural language data).
@@
Sparse-transformers [5] was one of the first works to introduce a sparse-attention method employing strided sparsity patterns in the attention mechanism. Reformer [17] used locally-sensitive hashing to compute attention scores in a sparse manner. Performer [6] used positive orthogonal random features to approximate the attention mechanism. Unlike these methods, which require training or fine-tuning, our approach operates entirely post-training without any fine-tuning.
@@
Another category of sparse attention methods employ token eviction policies to permanently delete tokens from the KV-cache based on some heuristic. StreamingLLM [35] uses initial tokens and a rolling KV-cache for processing infinite-length sequences. Zhang et al. [39] retain only "Heavy Hitters" tokens in the KV-cache based on accumulated attention scores. Scissorhands [20] prioritizes important tokens based on the "Persistence of Importance Hypothesis". Ge et al. [10] propose an adaptive eviction policy for each transformer layer. These methods are effective in reducing the memory and compute footprint of the attention but suffer from permanent loss of information leading to a non-trivial degradation in model quality. Our method does not involve any permanent loss of information with the trade-off of not reducing the memory footprint. Quantization-based approximate approaches [14, 23] are complementary to our work and can be applied in tandem.
@@
SparQ Attention [26] is a recent work that inspires our approach. They use high-magnitude query dimensions and corresponding key dimensions for approximate attention scoring, followed by computing the full attention scores for the top- keys. However, their method requires costly non-contiguous column indexing of the key vectors. Further, they store two copies of the past keys for efficiency, increasing memory use by 50%. In contrast, Loki avoids the extra memory and leverages the natural ordering of principal components, allowing for a more efficient slicing operation.
A concurrent work, InfiniGen [19], accelerates attention by pre-fetching top- keys from CPU to GPU memory, using SVD-based low-rank approximation of the attention scores. While their low-rank approximation is similar to Loki, our work provides deeper analysis of the intrinsic low-rank structure of attention keys and focuses on speeding up attention computation without CPU offloading. Importantly, their results affirm the benefits of the low-dimensional nature of attention keys applied in other contexts.
@@ 3 Dimensionality Analysis of Attention Keys
As noted in Section 1, Loki, our proposed method for sparse self-attention, is based on the observation that key tensors consistently reside in a lower-dimensional space than the full attention head dimension suggests. Here, we present empirical evidence supporting this claim by performing PCA on the keys generated in several language models and datasets.
3.1 Models and Datasets Used
To investigate the dimensionality of attention keys, we run 11 transformer-based models: Llama-2 7B/13B/70B [31], Llama-3 8B/70B [8], TinyLlama-1.1B [38], Pythia-6.9B [4], Mistral-7B [15], Mixtral-8x7B/8x22B [16], and Phi3-Mini-4K [22] on three popular English language datasets: WikiText-2 [21] (Validation Split), C4 [25] (Custom Split), and BookCorpus [40] (Custom Split). Custom splits are used for datasets where the validation split is not available. We run perplexity evaluation on these datasets and save the generated attention keys, before and after the application of rotary embeddings [29], referred to as pre-rotary and post-rotary keys, respectively throughout the paper. We then perform PCA on all the keys generated for each layer and head individually.
@@ 3.1 Models and Datasets Used
where, is the normalized eigenvalue of the covariance matrix of the keys for layer, and head, . We average this metric ranks across all heads of layer, and refer to it as .
3.2 Findings and Discussion
Figure 1 (left) shows the average across all layers for models with full key dimensionality of 128. We can see that the average rank is significantly lower than the full dimensionality of the keys for all models. Diving deeper, we present a layer-wise analysis for a few models: Llama2-7B, Llama3-70B, Mixtral-8x7B, and Phi3-Mini-4K in Figure 2. The results for the other models are similar and can be found in Appendix A.1.
We observe that the dimensionality of the keys (both pre-rotary and post-rotary) is significantly lower than the full dimensionality of the keys across all calibration datasets. Furthermore, the for a particular layer is consistent across datasets, for all combinations of models and datasets. This indicates that the lower-dimensional structure of the keys is consistent when calculated using different calibration datasets. Another trend we observe is that the initial layers of most models have a very low rank, as compared to the later layers, and this trend is particularly prominent for the pre-rotary keys. Lastly, we also observe that for most models, the average of across all layers is lower for pre-rotary keys as compared to post-rotary keys, indicating that the rotary embeddings increase the dimensionality of the keys. Further analysis on the variation of the rank across different heads within a layer and across different layers within a model can be found in Appendix A.1.
@@ 3.2 Findings and Discussion
4 Loki: Low-Dimensional Key Representations
We now describe our proposed algorithm for sparse attention – Loki. Loki leverages low dimensional projections of the keys in the KV-cache to efficiently and accurately select the top- (12.5-25%) most relevant tokens for self attention. Before discussing our approach, let us first look at some theoretical properties of attention in the PCA-transformed space of the key tensors.
4.1 Properties of Attention in the PCA-transformed Space
We begin by proving two lemmas that provide the rationale for our approach to compute attention in the PCA-transformed space.
@@ 4.1 Properties of Attention in the PCA-transformed Space
Let be the dimension of an attention head and be the PCA projection matrix of key tensors calibrated offline on a dataset. Assuming we are generating the token in the sequence, let be the query vector for the token, be the key vectors, including the past keys and the current key. Then, the attention scores computed using the PCA-transformed query and keys are equivalent to the attention scores computed using the original query and keys.
@@
Let be an orthogonal transformation that transforms the keys into the reduced dimension space as . Our objective is to minimize the following expression:
@@
We change our objective to minimize the upper bound in the RHS instead of the original objective. We know that PCA minimizes the reconstruction error (2nd term in the RHS) among all the orthogonal transformations. Thus, it follows that the optimal value of , and ∎
@@
Based on these lemmas and the inherent low-dimensional nature of key tensors in attention, we now introduce the workings of the Loki algorithm.
4.2 PCA-based Top-K Algorithm
Loki implements a PCA-based Top-K Attention approach. Previous works have shown that attention scores for a query are highly concentrated on a small subset of keys [36, 30]. This observation has motivated several methods that compute attention using only the top- most relevant keys. However, these previous works either compute the exact attention scores and then select the top- keys [12] or compute non-exact scores but have significantly higher memory requirements [26]. Loki alleviates these issues by computing approximate attention scores (for ranking the keys) in the reduced lower-dimensional space, without any significant increase in memory requirements. Algorithm 1 shows our Loki method. Line 5 of the algorithm computes the approximate attention scores using principal dimensions of the query and key vectors. Lines 6-7 select the top- keys based on the approximate attention scores. Line 8 computes the exact attention scores using the selected top- keys, directly in the transformed space (Lemma 4.1).
Compute and Memory Analysis: For vanilla attention, the complexity of computing is and the complexity of multiplying the values with the attention scores is . For Loki, the complexity of calculating the approximate attention scores (Line 5) is . The complexity of selecting the top- keys (Lines 6-7) is approximately (sorting followed by selection). The complexity of calculating the exact attention scores and multiplying with the values (Line 8-9) is . Additionally, the complexity of projections into the PCA space (Line 3) is . Assuming the complexity of selecting the top- keys is small compared to the other operations, the overall complexity of the algorithm is . Then, we have:
@@ 4.3 Implementation in Triton
Performing Loki efficiently involves complex indexing operations within the KV-cache (lines 5 and 7 of Algorithm 1). Standard PyTorch operations create temporary, dense copies of the KV-cache data in memory, leading to slowdowns due to expensive memory access. To alleviate this issue, we develop optimized kernels in Triton [1] for the three matrix multiplication operations in Loki. Our kernels can directly access relevant subsets of the KV-cache (both feature and sequence dimensions) and perform computations within GPU registers. This eliminates the need for creating dense copies, significantly improving performance. Our approach builds on SparQ [26], which introduced similar kernels for top- attention calculations. However, we identified and addressed inefficiencies in the SparQ kernels, which resulted in speedups of nearly in certain scenarios. (see Appendix C).
5 Experimental Setup
We evaluate Loki on the basis of perplexity using the WikiText-2 [21] dataset (test split), and on the basis of downstream task performance for short contexts using the LM-harness benchmark [9] and long contexts using LongBench [2]. For the short-context evaluation, we choose the same tasks and associated metrics as the HuggingFace OpenLLM leaderboard [3]. For the LongBench tasks, we evaluate on all the English language tasks.
We compare our method against three methods – full attention without any approximations, the exact TopK approach which computes the exact attention scores and then uses the top- tokens to compute the final output, and H2O [39], a popular token-eviction method. For these comparisons, we show the results with a budget size of = 0.25 and 0.125. For our method, we additionally use = 0.25 and 0.125. This configuration of our represents a 2.6x theoretical speedup. Table 1 provides an overview of the methods compared and the associated budget terms. H2O’s budget was split equally between the heavy hitter and recent tokens, as per the author’s recommendations. For H2O, we were unable to run the GSM8K task as the the author’s ML benchmarking code was too memory intensive to run for that task. For the aforementioned experiments, we generate PCA transforms using the WikiText-103 dataset. For the LongBench tasks, we compare our method with the full attention baseline as we were unable to run H2O due to memory constraints.
For the generalizability study, we compare the results of our method with PCA transforms from different calibration datasets: WikiText-103 [21], C4 [25], and BookCorpus [40]. Additionally, we also benchmark our triton based implementation of Loki by running an attention microbenchmark on a Llama2-13B-like setup (same hidden size and number of heads) for various prompt and generation lengths, and demonstrate speedups over vanilla attention.